OpenStudy (anonymous):

Math Analysis: Write an explicit formula for the nth term of the sequence. 18. 1, 1/3, 1/9, 1/27, 1/81

OpenStudy (amistre64):

well, 3,9,27,81 look familiar to me

OpenStudy (amistre64):

the 18. is a little off tho

OpenStudy (anonymous):

no, this is the equation

OpenStudy (amistre64):

do you know what a geometric sequence is?

OpenStudy (anonymous):

no, i forgot lol

OpenStudy (amistre64):

your times tables are geometric seqs. it is a sequence of numbers that are generated by multiply the same value to each term to get the next term

OpenStudy (anonymous):

Tn=ar^n-1

OpenStudy (anonymous):

that's the formula to find the indicated term of the geometric sequence i believe o-o

OpenStudy (amistre64):

hmm, i might be confusing the issue in my head your right about that being the general set up for a geo tho this might be an arithmatic sequence

OpenStudy (amistre64):

no, it geo ... im just getting too old lol

OpenStudy (amistre64):

3*3 = 9 9*3 = 27 27*3 = 81

OpenStudy (amistre64):

so, in your set up, a is the first term; and r is your common multiplier; or common ratio

OpenStudy (anonymous):

common ratio is 1/3?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

Tn=1 x 1/3^n-1

OpenStudy (amistre64):

redundant, but yes

OpenStudy (anonymous):

the answer at the back says tn=3(1/3)^n

OpenStudy (amistre64):

same thing

OpenStudy (anonymous):

how?

OpenStudy (amistre64):

youd have to remember rules for exponents to see the equalness about them

OpenStudy (amistre64):

a^(m-n) = a^m * a^n a^(-n) = 1/a^n

OpenStudy (anonymous):

so my answer is correct? if i put that answer down on the test i won't get marked down cause this is for my final lol

OpenStudy (amistre64):

1/3^(n-1) = 1/3^n * 1/3^(-1) = 1/3^n * 3

OpenStudy (amistre64):

im not the one grading it

OpenStudy (anonymous):

so my answer is wrong?

OpenStudy (amistre64):

your answer is equivalent; but how your spose to get it to LOOK is up to the grader

OpenStudy (amistre64):

some people dont wanna see negative exponents; others could care less

OpenStudy (anonymous):

my teacher wants us to simplify our answers

OpenStudy (anonymous):

idk if my answer was simplified

OpenStudy (amistre64):

the 1x part is redundant, and can be simplified by ignoring it. 1x anything doesnt change its value

OpenStudy (amistre64):

\[T_n=(\frac{1}{3})^{n-1}\] rewrite it however you see fit

OpenStudy (anonymous):

so now that answer is simplified?

OpenStudy (amistre64):

.... there is no standard for "simplified". It is solely up to the person who is grading and what they have taught you

OpenStudy (amistre64):

since i was not there, im wouldnt know

OpenStudy (anonymous):

okay, but i think the answer on my answer sheet would be preferred by my teacher o-o

OpenStudy (amistre64):

it should be :)

OpenStudy (anonymous):

can you show me a way how to get tn=3(1/3)^n

OpenStudy (amistre64):

i did, its posted above

OpenStudy (anonymous):

what if i use this equation? an= a + (n-1) d

OpenStudy (amistre64):

use it for what?

OpenStudy (anonymous):

for #18

OpenStudy (amistre64):

since there is not "d" (common difference) between the terms; it would amount to trying to place a square peg into a round hole

OpenStudy (anonymous):

so i use a different equation?

OpenStudy (amistre64):

we have already used the appropriate equation; nothing else will work for it.

OpenStudy (anonymous):

a^(m-n) = a^m * a^n a^(-n) = 1/a^n

OpenStudy (anonymous):

this one rite?

OpenStudy (amistre64):

those are rules that govern exponents and how to rewrite this to fit your books answer

OpenStudy (anonymous):

oh, so that was not an equation?

OpenStudy (amistre64):

no, they werent. those are generalized rules for exponents that show you what you can do to them

OpenStudy (anonymous):

ok, so why does 1/3^-1 turn into 3?

OpenStudy (anonymous):

ohh, i get it haha :)

OpenStudy (anonymous):

i use my calculator, but for this part i can't use my calculator

OpenStudy (amistre64):

a negative exponent is a denominator is the simplest explanation i can think of

OpenStudy (anonymous):

theres one part where i cant use a calculator for the test

OpenStudy (anonymous):

so, if a number is to the -1 power u just put the number on the bottom to the top?

OpenStudy (anonymous):

like 1/3

OpenStudy (amistre64):

correct; we just flip the number

OpenStudy (anonymous):

okay ty :)

OpenStudy (amistre64):

1/3 ^ -1 flips to 3 yw